Can someone explain this to me please help ASAP!!!!!!!!!!!!!
During the last hour, a telemarketer dialed 25 numbers and reached 7 busy signals, 3 answering machines, and 15 people. Use this information to determine the empirical probability that the next call will be answered in person

thank you

Answer by checkley77(12844)

15/25=3/5=.6=60% CHANCE OF GETTING A PERSON ON THE NEXT CALL.